### v5

# S3E: A Large-scale Multimodal Dataset for Collaborative SLAM

Abstract
The burgeoning demand for collaborative robotic systems to execute complex tasks collectively has intensified the research community’s focus on advancing simultaneous localization and mapping (SLAM) in a cooperative context. Despite this interest, the scalability and diversity of existing datasets for collaborative trajectories remain limited, especially in scenarios with constrained perspectives where the generalization capabilities of Collaborative SLAM (C-SLAM) are critical for the feasibility of multi-agent missions. Addressing this gap, we introduce S3E, an expansive multimodal dataset. Captured by a fleet of unmanned ground vehicles traversing four distinct collaborative trajectory paradigms, S3E encompasses 13 outdoor and 5 indoor sequences. These sequences feature meticulously synchronized and spatially calibrated data streams, including 360-degree LiDAR point cloud, high-resolution stereo imagery, high-frequency inertial measurement units (IMU), and Ultra-wideband (UWB) relative observations. Our dataset not only surpasses previous efforts in scale, scene diversity, and data intricacy but also provides a thorough analysis and benchmarks for both collaborative and individual SLAM methodologies. For access to the dataset and the latest information, please visit our repository at https://pengyu-team.github.io/S3E.
Index Terms:
Collaborative Simultaneous Localization and Mapping (C-SLAM), Multi-Agent Systems, Dataset, Trajectory Paradigms
I Introduction
Collaborative Simultaneous Localization and Mapping (C-SLAM) is pivotal for enabling multi-agent cooperation, thereby enhancing the robustness and efficiency of concurrent localization and mapping endeavors within a shared operational space [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Despite considerable progress in C-SLAM research during the past decade, issues of repeatability and benchmarking remain significant hurdles within the domain. The field is primarily constrained by two limitations. First, the complexity of C-SLAM systems, which integrate intricate software architectures and a variety of sensors, complicates the process of achieving perfect replication, thereby emphasizing the need for publicly available datasets to facilitate methodological evaluation. Second, a common method for evaluating C-SLAM solutions involves partitioning a single-agent SLAM dataset into several segments, each assigned to a distinct virtual agent [11]. However, this methodology is impractical for real-world multi-agent scenarios, given the uniform perspectives and lighting conditions that all virtual agents would encounter at intersection points.
Furthermore, the trajectory design for multi-robot scenarios must adhere to two fundamental principles. The first principle is temporal and spatial diversity. The trajectories should be designed to accommodate the varying intra- and inter-loop closures at different times and locations, thus offering complementary vantage points for the participating robots. The second principle pertains to the communication constraints. In practical applications, collaborative robots are limited to sharing information within a certain proximity. Hence, maintaining a reasonable interaction distance between robots within these regions is essential. In this paper, we introduce four trajectory prototypes that conform to diverse intra- and inter-robot closure strategies, aiming to evaluate the generalizability of C-SLAM methodologies.
Within the SLAM research community, publicly available datasets are indispensable for accelerating the development of SLAM algorithms suitable for real-world deployment. This significance stems from two primary factors. First, the execution of specialized SLAM experiments requires substantial investment in sophisticated hardware and complex software, including calibration routines and ground truth generation systems; hence, the availability of datasets substantially aids in the algorithmic development process. Second, these datasets provide a fair and impartial benchmarking framework for comparative analysis and performance assessment. Over the past decade, an array of SLAM datasets has emerged, each contributing uniquely to the diversity of sensor modalities and the breadth of scenarios and motion types they encompass [12, 13, 14, 15, 16, 17]. Nonetheless, the preponderance of these datasets is skewed towards single-agent systems, equipped with an assortment of sensors such as cameras, inertial measurement units (IMUs), and LiDAR, which may limit their applicability to multi-agent contexts.
The limited availability of datasets specifically tailored for C-SLAM is largely due to the escalating challenges in data acquisition as the number of collaborating agents increases. Existing C-SLAM datasets, such as those from [18, 19, 20, 21, 22, 23], predominantly feature a single cooperative paradigm with substantial spatial overlap among the robots, often overlooking cooperative strategies in areas with minimal overlap. To address this gap and bolster ongoing research in C-SLAM, we introduce S3E, a comprehensive multimodal dataset. S3E captures a variety of cooperative trajectory patterns across both outdoor and indoor settings.
Utilizing our S3E dataset as a foundation, we have conducted an extensive series of experiments employing state-of-the-art SLAM methodologies. These experiments spanned the spectrum of SLAM applications, from single-robot to collaborative scenarios, incorporating a rich array of sensor data. Our results indicate that current collaborative SLAM techniques often encounter suboptimal performance, particularly in handling inter-loop closures across certain sequences. This finding highlights the critical need for ongoing research aimed at improving the robustness and reliability of C-SLAM systems.
In conclusion, our work makes several key contributions to the field:
  • We have assembled a pioneering C-SLAM dataset through the deployment of three state-of-the-art ground robots. Each robot is equipped with an advanced 16-beam 3D laser scanner, two high-resolution color cameras, a 9-axis IMU, a UWB receiver, and a dual-antenna RTK receiver. Notably, this is the first C-SLAM dataset to include UWB relative distance measurements, offering a novel dimension for research.
  • To evaluate the adaptability of C-SLAM in environments with sparse overlap, we have captured extensive long-term sequences. These were obtained using four carefully crafted trajectory paradigms that account for a variety of intra- and inter-robot loop closure scenarios.
  • We have undertaken an in-depth evaluation of contemporary C-SLAM techniques and their single-agent counterparts. This assessment leverages a wide array of sensor data, with a meticulous focus on the unique characteristics and performance attributes of each method.
[…]
[autoplay,loop,width=]5videos/360/dormitory_1/000119
Figure 1: Animations of Dormitory_1. Data is captured simultaneously by a trio of Unmanned Ground Vehicles (UGVs). Each UGV initiates its journey from distinct starting points and independently explores separate areas. Upon successful detection of inter-loop closures, they exchange valuable information and collaboratively construct a globally consistent map. For optimal viewing of the animations, it is recommended to utilize Adobe Reader or KDE Okular.
Table I: Comparative Analysis of Prominent SLAM Datasets. The abbreviations used within the table are as follows: "Sw" denotes Software Synchronization, indicating that synchronization across sensors or systems is achieved through software mechanisms; "Hw" represents Hardware Synchronization, where synchronization is managed by hardware means.
Dataset Platform Sensors Time Sync. Trajectory Environment Ground Truth
Camera IMU LiDAR UWB Intra- Inter- Overlap Paradigm Indoor Outdoor
KITTI [12] Car Sw - - GNSS/INS
EuRoC [13] UAV Hw - - Motion Capture
UTIAS [18] 5 UGVs Sw NTP Large 1 Motion Capture
AirMuseum [20] 3 UGVs, 1UAV Sw NTP Large 1 SfM
FordAV [19] 3 Cars - GNSS Large 1 GNSS/INS
GRACO [22] 1UGV, 1UAV Hw - Restricted 1 GNSS/INS
Kimera-Multi [21] 8 UGVs - - Restricted 1 GPS and total-station assisted LiDAR SLAM
SubT-MRS [23] UGVs, UAVs, Handheld Hw - Restricted 1 3D Scanner
Ours 3 UGVs Hw GNSS, PTPv2 Restricted 4 RTK, GNSS/INS Motion Capture
(a) Sensor locations and the coordinate frames.
(b) S3Ev1.0.
(c) S3Ev2.0.
Figure 2: Mobile Platform Sensor Layout and Coordinate Systems. The left part details the sensor locations and the coordinate frames that define their spatial orientation relative to the platform. In the right part, our mobile platforms are available in two versions, each designed for different operational requirements.
II Related Work
Collaborative SLAM. The integration of collaborative agents within the framework of SLAM presents a significant advancement, offering the potential for increased operational velocity and efficiency. This collaborative approach has the capacity to outstrip the performance of single-agent systems [6, 24]. LABEL:Fig:ani_dorm_1 demonstrates the operational concept of C-SLAM, where a collective of robots is enabled to explore various sectors of an environment in parallel, thereby enhancing the overall efficiency of the exploration process compared to traditional single-agent methodologies.
C-SLAM systems can be differentiated based on their underlying network topologies, which include centralized, distributed, and adaptive models. In a centralized topology, a central server is tasked with the aggregation of all incoming sensory data, ensuring a unified estimation of multiple trajectories, and subsequently distributing these findings to the constituent agents [25, 26, 4]. Conversely, in distributed topologies, computational responsibilities are decentralized among the team of agents, with some systems utilizing a message broadcasting protocol for the propagation of information [27]. However, scalability becomes a formidable challenge in such distributed frameworks, particularly in scenarios involving a large number of robots, due to limitations in bandwidth and communication range.
To mitigate these challenges, adaptive strategies have been proposed, advocating for the transmission of information within the immediate communication sphere of the agents [28, 1, 5]. Swarm-SLAM [29] represents an innovative paradigm within this domain, adept at minimizing the communication load between robots while simultaneously accelerating the convergence of the SLAM process through the implementation of sophisticated prioritization algorithms for inter-robot loop closures. Moreover, hybrid approaches present a viable alternative, wherein a subset of robots engages in data processing, while others are dedicated to the application of the outcomes derived from this processing [30].
SLAM Datasets. High-quality and extensive datasets are indispensable for propelling advancements in SLAM research. Over recent years, there has been a commendable effort to release datasets to the research community. The KITTI dataset [12] is particularly noteworthy for its extensive and high-quality urban data, serving as a benchmark for both vision-based and LiDAR-related SLAM techniques. However, KITTI is not without limitations; it lacks perfectly synchronized data, offering only estimated time delays between different sensors. To surmount this limitation, the EuRoC dataset [13] was collected, providing synchronized visual and inertial data captured by a micro aerial vehicle in controlled environments, synchronized through the use of hardware clocks.
Despite these commendable efforts, there is a recognized scarcity of datasets tailored for multi-agent scenarios. The UTIAS dataset [18] stands as a pioneering work in multi-agent SLAM, featuring a collaborative system of five robots exploring a small area with identifiable landmarks. The AirMuseum dataset [20] introduces a heterogeneous multi-robot SLAM dataset within a larger warehouse environment, offering ground-truth trajectories computed using Structure-from-Motion (SfM) [31] and constrained with fixed beacons. The FordAV dataset [19] presents a dataset of multiple autonomous vehicles with a 3D map, capturing seasonal variations in dynamic urban environments. However, these datasets tend to have substantial overlap in the co-view region, which may limit the comprehensive evaluation of C-SLAM robustness.
In an effort to address the limitations of existing datasets, the GRACO dataset [22] presents a novel approach by introducing a ground-aerial dataset. This dataset leverages the drone’s altitude to dynamically adjust the extent of view overlap, offering a unique perspective on multi-agent SLAM scenarios. However, GRACO exhibits some inconsistencies in the synchronization of data acquisition between aerial and ground perspectives, and it does not fully accommodate the complexities of dynamic environments.
The SubT-MRS dataset [23] contributes to the field by providing a diverse array of multi-robot SLAM data captured across a variety of environments. This dataset enhances the representation of real-world applications and the challenges posed by different settings.
Kimera-Multi [21] further advances the state of C-SLAM datasets by offering a comprehensive collection characterized by its diverse trajectory designs and the involvement of multiple robots. This dataset is instrumental in studying the collaborative aspects of multi-agent systems.
It is crucial to acknowledge, however, that existing datasets, including GRACO, SubT-MRS, and Kimera-Multi, often do not fully consider the intricacies of collaborative trajectory patterns. The robustness of C-SLAM systems is influenced by the ability to handle intra- and inter-loop closures, a factor that is not consistently addressed in current datasets. The consideration of these collaborative patterns is essential for developing C-SLAM algorithms that can effectively operate in environments with varying levels of agent interaction and dynamic scene elements.
In this paper, we introduce a large-scale C-SLAM dataset that includes visual ambiguities and dynamic objects, captured under four distinct trajectory paradigms to assess C-SLAM generalizability under restricted overlap conditions. Our dataset also incorporates UWB relative measurement positioning, adding a significant dimension for C-SLAM research.
For a comprehensive comparison of different datasets, including our contribution, refer to LABEL:Tab:dataset_comparison.
Table II: Payload Sensor and Ground Truth Device Specifications. It includes information about the types of sensors, their respective capabilities, and the ground truth devices used for validation and accuracy assessment.
Device Type Unit
LiDAR Velodyne
   Model VLP-16 Puck
   Channels $16$
   Range $100$ $m$
   Accuracy ±$\pm 3$ $cm$
   Vertical FOV ±$\pm 15$
   Vertical Resolution $2$
   Horizontal FOV $360$
   Horizontal Resolution $0.1\sim 0.4$
   Max. Frequency $10$ $Hz$
Camera - Stereo HikRobot
   Model MV-CS050-10GC
   Data Interface GigE
   Sensor Model Sony IMX264
   Resolution $1224\times 1024$ $pixel$
   Dynamic Range $72$ $dB$
   Baseline $360$ $mm$
   Max. Frequency $24.2$ $Hz$
   Lens Model MVL-MF0828M-8MP
   Focal Length 8 $mm$
   Vertical FOV 54.97
   Horizontal FOV 68.46
IMU Xsens
   Model MTi-30-2A8G4
   Gyro In-Run Bias Stability $18$ ${}^{\circ}/h$
   Accel In-Run Bias Stability $15$ $\mu g$
   Mag RMS Noise 0.5 $mGauss$
   Max. Frequency 400 $Hz$
UWB NoopLoop
   Model LinkTrack P-B
   Max. Distance $500$ $m$
   Accuracy ±$\pm 10$ $cm$
   Max. Frequency $200$ $Hz$
RTK Femtomes
   Model Nano-D
   RTK Accuracy ±$\pm 1$ $cm$
   Max. Frequency 20 $Hz$
GNSS/INS CHCNAV
   Model CGI-610
   RTK Accuracy ±$\pm 1$ $cm$
   Gyro In-Run Bias Stability 2.5 ${}^{\circ}/h$
   Accel In-Run Bias Stability 15 $\mu g$
   Max. RTK Frequency 20 $Hz$
   Max. INS Frequency 100 $Hz$
Motion Capture Vicon
   Model Tracker 3.8
   Camera Model Vero 2.2
   Num. of Cameras 17
   Position Accuracy ±$\pm 1$ $mm$
III S3E Dataset
The S3E dataset has been meticulously assembled, prioritizing high temporal precision in the synchronization of data from an array of sensors. Each sensor is calibrated to a shared timescale, ensuring that the captured multi-sensory information is harmonized. This precise synchronization is facilitated by an advanced time synchronization mechanism, essential for the accurate integration and co-registration of data across diverse modalities. For a visual representation of the sensor placement on our mobile robot platform, refer to LABEL:Fig:sensor_layout. Our mobile platforms are offered in two specialized versions, each tailored to meet specific operational needs:
  • S3Ev1.0 (LABEL:Fig:platform_v1.0): Customized for indoor operations, this version boasts a compact design that guarantees exceptional maneuverability, enabling it to traverse tight indoor spaces effortlessly.
  • S3Ev2.0 (LABEL:Fig:platform_v2.0): This enhanced version features an expanded width, accommodating a state-of-the-art, higher-precision ground truth system. The integration of an UWB module significantly bolsters the platform’s localization accuracy, particularly in environments where traditional SLAM systems may falter.
The subsequent sections delve into a comprehensive description of the S3E dataset’s components and their distinctive features, highlighting the meticulous approach taken to ensure the dataset’s utility and reliability in advancing SLAM research.
III-A Sensor Configuration
(a) S3Ev1.0.
(b) S3Ev2.0.
Figure 3: Time Synchronization Scheme. The scheme addresses synchronization in both outdoor environments using GNSS time as the global reference and in GNSS-denied environments like indoors and tunnels, where agents rely on a PTPv2 server for time synchronization.
Our S3E dataset encompasses a multimodal array of sensors, each selected for its operational range and noise characteristics, and meticulously synchronized to capture data with high temporal precision. The sensors are integrated onto the Agilex Scout Mini, a versatile all-terrain mobile platform capable of high-speed remote-controlled navigation, featuring four-wheel drive and a maximum speed of $10km/h$.
For a comprehensive understanding of the data acquisition process and to facilitate replicability of our experiments, LABEL:Tab:hardware_properites outlines the technical specifications of the sensors and ground truth devices that were integrated into our payload. This includes the sensor types, their resolution, measurement range, accuracy, and any other pertinent technical details that define their contribution to the SLAM system’s performance.
III-B Sensor Synchronization
This section delves into the critical processes of time synchronization and sensor calibration, which are essential for achieving optimal sensor fusion and maximizing system performance within a dynamic, multimodal environment.
III-B1 Time Synchronization.
Time synchronization across all sensors is essential for the accurate co-registration of multi-sensor data, particularly in dynamic environments where precise timing is critical to the quality of the fused dataset. This synchronization ensures that data from various sensors is temporally aligned, facilitating coherent integration and analysis.
Our synchronization system is built around an Altera EP4CE10 FPGA board acting as the primary trigger device, with an Intel NUC11TNKv7 serving as the host computer. The system is designed with a comprehensive set of I/O interfaces to accommodate the diverse requirements of different sensors, as depicted in LABEL:Fig:Time-Synchronization-Scheme..
For synchronization across agents, we address two distinct scenarios:
  • In outdoor settings with access to Global Navigation Satellite System (GNSS) signals, we use GNSS time as the global reference to synchronize the timing across agents.
  • In GNSS-denied environments such as indoors and tunnels, agents synchronize their timers by obtaining external global time data from Alpha, which serves as a Precision Time Protocol version 2 (PTPv2) server, via a wireless network connection.
Within each agent, the FPGA orchestrates the timing for various sensors:
  • S3Ev1.0: The FPGA periodically issues a pulse from a low-drift oscillator to trigger the LiDAR, stereo cameras, and IMU, as shown in LABEL:Fig:time_sync_v1.0. Specifically, the FPGA generates a 1 Hz pulse to initiate the LiDAR, which then outputs data at 10 Hz and updates its internal counter register upon receiving the trigger signal. The stereo cameras and IMU are designed to immediately return data following the receipt of an FPGA trigger pulse. The FPGA is configured to generate trigger pulses at 10 Hz for the cameras and 100 Hz for the IMU, tailored to the specific data output needs of these sensors.
  • S3Ev2.0: This version enhances synchronization by aligning the FPGA board with the pulse-per-second (PPS) signal from its internal GNSS/INS module, as depicted in LABEL:Fig:time_sync_v2.0. This alignment generates frequency trigger signals for IMU, stereo cameras, and LiDAR clock synchronization. The UWB receiver, not supporting trigger-based data acquisition, collects and outputs measured distance data directly.
Considering transmission delays, all sensor readings are forwarded to the host computer, where they are timestamped upon arrival, organized, and packaged to ensure accurate temporal referencing for subsequent analysis.
III-B2 Sensor Calibration.
Our sensor suite is governed by a unified coordinate system framework, adhering to the right-hand rule, which ensures uniformity in data orientation and facilitates standardized analysis.
The calibration process was meticulously carried out in a series of five distinct stages, each designed to ensure the precise alignment and integration of sensor data:
Camera Intrinsic Calibration: We executed a high-precision intrinsic calibration for our cameras using the conventional checkerboard pattern, a method renowned for its reliability [32]. This process determined the camera intrinsic matrix $K$ and the lens distortion coefficients $D$, which are essential for correcting lens distortions and enhancing image accuracy.
IMU Intrinsic Calibration: Intrinsic calibration of the IMU involved accounting for white noise $\sigma_{n}$ and random walk noise $\sigma_{m}$ in the measurements. Building upon the nominal values provided in the Xsens MTi-30-2A8G4 IMU’s data sheet, we refined these parameters through a calibration process utilizing Allan variance, thereby achieving a more accurate probabilistic model of IMU noise [33].
Camera-Camera Extrinsic Calibration: The extrinsic calibration of the stereo cameras was performed to determine the rotation matrix $R$ and the translation vector $t$, which were computed by minimizing reprojection errors. This step is crucial for establishing the geometric relationship between the two cameras in a stereo setup [32].
Camera-IMU Extrinsic Calibration: The spatial transformation $T_{ic}$ between the camera and IMU was ascertained using the open-source Kalibr toolbox [34]. This calibration is essential for the integration of IMU data with visual data for enhanced pose estimation and sensor fusion.
Camera-LiDAR Extrinsic Calibration: The extrinsic calibration between the camera and LiDAR, represented as $T_{lc}$, was established by minimizing line-to-plane errors, ensuring the accurate co-registration of LiDAR point clouds with camera imagery [35].
Upon completion of these calibration stages, the resulting parameters were meticulously stored in YAML files for each agent, ensuring that the calibrated parameters are readily accessible and reproducible.
III-C Ground Truth
To establish accurate ground truth tracks for our dataset, we implement a multifaceted approach leveraging three distinct methods.
In outdoor settings where GNSS signals are readily accessible, we employ a dual-antenna Real-Time Kinematic (RTK) device to capture localization data with an accuracy reaching centimeter-level precision. For the S3Ev2.0 platform, we have made a significant advancement in our ground truth instrumentation by integrating a GNSS/INS unit. This upgrade enables high-frequency positioning data output, even during GNSS signal outages, such as those encountered in tunnel environments. The system achieves this by fusing RTK measurements with data from the built-in IMU through a post-processing fusion technique, thereby ensuring continuous and reliable ground truth data for our dataset.
In GNSS-denied environments, such as indoor laboratories, we employ a motion capture system equipped with 17 high-frequency cameras to record the start and endpoints of the tracks. This setup is utilized due to the prohibitive costs associated with constructing a large-scale motion capture system. For indoor scenarios not equipped with a motion capture system, such as in teaching buildings, we record the initial and terminal points of the tracks using the RTK device.
By combining these methods, we are able to generate comprehensive ground truth positions that cater to both outdoor and indoor environments, thereby facilitating a robust evaluation of C-SLAM algorithms across diverse real-world scenarios.
Figure 4: Illustration of Trajectory Paradigms for C-SLAM. The distinct trajectory paradigms adopted by three agents, designated as Alpha, Bob, and Carol, to demonstrate different interaction and information exchange patterns in a multi-agent SLAM context. (a) Concentric Circles. (b) Intersecting Circles. (c) Intersection Curve. (d) Rays
III-D Trajectory Paradigms
The S3E dataset incorporates a diverse set of trajectory paradigms to simulate various collaborative scenarios that agents may encounter during SLAM operations. As despited in LABEL:Fig:trajecotry_paradigms, we have designed these paradigms to address different loop closure conditions, both within individual agents (intra-loop) and across multiple agents (inter-loop). The following are the four trajectory paradigms we have included in our dataset:
Concentric Circles: In this paradigm, multiple unmanned platforms move in a coordinated formation, maintaining interconnectivity and observing consistent environmental data. This setup enables continuous loop detection among plat- forms, enhancing the system’s ability to perform comprehensive area coverage and exploration tasks such as environmental monitoring and map construction. The real-time sharing of perceptual information and pose estimation results among platforms bolsters the system’s positioning and mapping accuracy. However, the high similarity of observed data may complicate the detection and management of dynamic environmental changes.
Intersecting Circles: This paradigm is characterized by individual unmanned platforms operating in distinct zones, exchanging perceptual information upon intersection. Communication is established only at these points, with partial environmental information overlap. The exchange of historical data for loop closure detection at intersections, along with independent loop closure capabilities of each platform, expands the exploration range and improves mission efficiency. The division of search areas reduces communication and coordination demands within the system. However, limited interaction between platforms may hinder the achievement of globally consistent mapping and positioning.
[…]
Trajectory Paradigm Closures Applications
Intra-loop Inter-loop
[…]
Intersection Curve: Unmanned platforms initiate from diverse starting points, intersect at designated locations, and converge at a common endpoint. Communication and loop closure detection occur only upon meeting, with partial environmental information overlap. This paradigm is adept for large-area distributed exploration and mapping, facilitating regular information exchange and loop closure detection to minimize cumulative errors. The independent exploration by each platform enhances task efficiency and coverage. However, higher navigation and control capabilities are required for planning suitable intersection points and paths.
Rays: In this paradigm, platforms depart from various starting points, moving independently until converging at a shared endpoint. Communication is restricted to the final convergence, with minimal environmental information overlap. The platforms can exchange historical information for loop closure detection at convergence, but lack individual loop closure detection capabilities. This paradigm is suitable for large-scale independent exploration and mapping tasks, such as post-disaster searches and mine explorations. It emphasizes the autonomous execution of tasks by each platform, relying on its perceptual and computational resources. The final convergence allows for historical information exchange and local map optimization. However, the scarcity of information exchange poses challenges for globally consistent mapping and positioning, necessitating advanced autonomous navigation and decision-making.
Each paradigm is meticulously crafted to offer a robust framework for assessing C-SLAM algorithms across a multitude of real-world collaborative robotic applications. LABEL:Tab:Trajectory-Paradigms. provides a comparative analysis of the loop closure detection situations corresponding to these paradigms, highlighting their respective advantages and potential application areas.
Table IV: Analysis of S3E dataset.
Scenario Time[s] Trajectory Ground Truth Length[m] Size[GB] Sensors
Env. Region a b c d Alpha Bob Carol Camera IMU LiDAR UWB
Outdoor v1.0 Square_1 460 RTK 546.0 496.5 529.2 17.8
Square_2 255 RTK - 250.6 246.4 9.4
Library_1 454 RTK 507.6 517.2 498.9 16.3
Campus_Road_1 878 RTK 920.5 995.9 1072.3 29.4
Playground_1 298 RTK 407.7 425.6 445.5 8.7
Playground_2 222 RTK 265.6 315.7 456.4 6.3
Dormitory_1 671 RTK 727.0 719.3 721.9 23.5
v2.0 Square_3 466 GNSS/INS 487.4 569.6 563.8 8.5
Library_2 491 GNSS/INS 521.6 523.9 520.3 9.3
Campus_Road_2 1594 GNSS/INS 1938.6 1934.2 1950.1 30.9
Campus_Road_3 907 GNSS/INS 983.3 967.6 986.9 17.5
Playground_3 111 GNSS/INS 84.7 91.3 110.7 1.8
Tunnel_1 425 GNSS/INS 521.9 502.4 501.1 8.2
Indoor v1.0 Teaching_Building_1 798 RTK 617.2 734.4 643.4 27.3
Laboratory_1 292 Motion Capture 147.7 161.5 141.0 9.6
Laboratory_2 391 Motion Capture 215.3 199.1 160.3 12.7
Laboratory_3 410 Motion Capture 219.1 202.2 204.2 13.3
Laboratory_4 380 Motion Capture 173.7 177.2 180.0 12.7
[…]
.1 Refer to caption S3E. .2 Refer to caption <platform_id>. .3 Refer to caption Calibration. .4 Refer to caption <agent_id>.yaml. .3 Refer to caption S3E_<sequence_id>. .4 Refer to caption S3E_<sequence_id>.db3. .4 Refer to caption metadata.yaml. .4 Refer to caption groudtruth.txt.
Figure 6: S3E Dataset Organizational Structure.
III-E Dataset Analysis
To align with the collaborative mission scenarios and to reflect the nuances of the four trajectory paradigms, we have strategically chosen a variety of complex and representative scenes across different areas within Sun Yat-sen University’s Guangzhou East Campus for our data collection endeavors. The mobile platforms, comprising three tele-operated robots, are under human control to ensure safety during operations.
For an exhaustive overview of the trajectory paradigms applied, the lengths of the trajectories, and the types of sensors utilized for each sequence, please refer to LABEL:Tab:Analysis-of-S3E. The dataset encompasses a diverse range of environments, including squares, libraries, and six additional regions, comprising a total of 13 outdoor and 5 indoor sequences. The aggregated size of all 18 sequences is over 263.2 gigabytes, with a combined duration of more than 475.2 minutes and an extensive operating distance surpassing 28.1 kilometers. On average, the sequences last for 527.9 seconds each, which validates the dataset’s adequacy for extensive, long-term assessments of C-SLAM technologies.
Furthermore, the dataset encompasses a variety of trajectories that adhere to the paradigms outlined in Section LABEL:Subsec:Trajectories. This diversity is instrumental for rigorous testing and comparative analysis of different C-SLAM methodologies under various interactive conditions, such as intra-robot and inter-robot loop closures.
To thoroughly assess the accuracy and robustness of C-SLAM algorithms in complex, real-world scenarios, our dataset encompasses a diverse range of environments, each presenting unique challenges:
Dormitory: Characterized by high pedestrian traffic and the presence of dynamic objects such as pedestrians and bicycles, dormitory areas challenge the perception and tracking abilities of unmanned platforms. Their regular architectural layouts offer an ideal setting for evaluating the precision and consistency of C-SLAM algorithms across multiple platforms.
Campus Road: Serving as critical connectors, campus roads are marked by long distances and expansive views. They test the endurance and large-scale exploration capabilities of unmanned platforms. Extensive data collection along these roads, including long-distance and multi-cycle datasets, provides a robust foundation for assessing the stability, accuracy, and efficiency of C-SLAM algorithms over extended operations.
Playground: As open areas with fewer obstructions, sports fields challenge feature extraction, registration, and optimization processes. Data collected at various times of day, including day and night, evaluates the adaptability of C-SLAM algorithms to different lighting conditions. Additionally, data from rapid motion scenarios assesses the algorithms’ performance under significant motion.
Laboratory: Indoor environments like laboratories, with their confined spaces, complex layouts, and rich semantic content, challenge navigation, obstacle avoidance, and semantic mapping. The presence of various instruments and furniture makes laboratories suitable for testing the advanced scene understanding and mapping capabilities of C-SLAM algorithms.
Teaching Building and Tunnel: These areas, with their severe perceptual aliasing, pose significant challenges due to poor lighting in tunnels and similar geometric structures in corridors, which can lead to errors in data association. They test the robustness of C-SLAM algorithms in maintaining accurate positioning and mapping.
Furthermore, our dataset includes other distinctive regions such as libraries and squares, which contribute additional environmental features and perceptual challenges to the dataset’s diversity and comprehensiveness. LABEL:Fig:outdoor_tracks displays the various trajectory paradigms captured in outdoor environments within the S3E dataset. Our dataset encompasses a diverse array of challenging environments that are likely to be faced by C-SLAM algorithms in real-world deployments. These include dynamic settings with moving objects, prolonged operational durations, scenarios with perceptual aliasing, indoor environments, and situations characterized by substantial motion. The inclusion of such a wide range of conditions lays a solid groundwork for an in-depth evaluation of the performance metrics, adaptability to different conditions, and overall robustness of C-SLAM algorithms. This thorough vetting process is essential for advancing the state-of-the-art in collaborative robotic navigation and mapping technologies.
Table V: Descriptions of ROS2 topics included in our dataset. Note that agent_id stands for Alpha, Blob, and Carol.
Sensor Topic Type Description Frequency
LiDAR /agent_id/velodyne_points sensor_msgs/msg/PointCloud2 Raw velodyne pointcloud with $[x,y,z,intensity,ring,time]$ 10 $Hz$
Camera11footnotemark: 1 /agent_id/left_camera/compressed sensor_msgs/msg/CompressedImage Compressed RGB8 image with $1224\times 1024$ 10 $Hz$
/agent_id/right_camera/compressed sensor_msgs/msg/CompressedImage Compressed RGB8 image with $1224\times 1024$
IMU /agent_id/imu/data sensor_msgs/msg/Imu Raw imu data 100 $Hz$
RTK22footnotemark: 2 /agent_id/fix sensor_msgs/msg/NavSatFix RTK-GNSS data with $[latitude,longitude,altitude]$ 1 $Hz$
UWB33footnotemark: 3 /agent_id/nlink_linktrack_nodeframe2 std_msgs/msg/Float64MultiArray UWB data comprises 13 data fields 44footnotemark: 4 50 $Hz$
GNSS/INS33footnotemark: 3 /agent_id/time_reference sensor_msgs/msg/TimeReference GPS time reference anchored at the epoch of 1980-01-06 00:00:00 UTC 100 $Hz$
/agent_id/fix sensor_msgs/msg/NavSatFix RTK-GNSS data with $[latitude,longitude,altitude]$
/agent_id/correct_imu sensor_msgs/msg/Imu Raw build-in imu data
/agent_id/vel geometry_msgs/msg/TwistStamped Timestamped linear velocities $(m/s)$ of mobile platform along $[east,north,up]$
/agent_id/heading geometry_msgs/msg/QuaternionStamped Timestamped quaternions map mobile platform attitudes
  • 11footnotemark: 1
    Available in all sequences of S3Ev1.0. 22footnotemark: 2 Available in outdoor sequences of S3Ev1.0. 33footnotemark: 3 Available in all sequences of S3Ev2.0.
  • 44footnotemark: 4
    Detailed definitions of UWB data fields are documented and accessible on our project website.
III-F Dataset Format
We employ the ROS2 [36] bag format for storing the sensor data, which is a widely-recognized standard in the robotics community. This format allows for efficient data management and playback, enabling users to simulate and analyze the dataset in various software environments. For the purpose of simplifying data replay and ensuring seamless access to the synchronized sensor streams from our robotic platforms, we have implemented a consolidation strategy. Specifically, we merge the sensor data collected from the Alpha, Blob, and Carol robots during the same operational sequence into a single ROS2 bag file.
Data Organization: As depicted in LABEL:Fig:data_organization , our dataset features a well-organized directory structure. The calibration parameters are documented in YAML files and stored in the dedicated Calibration directory. Each sequence is accompanied by essential data files, including .db3 file that house the primary sensor measurements, and metadata.yaml file that provide essential information about the sequence, such as sensor settings and capture conditions. Furthermore, to facilitate comprehensive performance evaluation, we have included auxiliary files such as groundtruth.txt, which contains the ground truth data.
Ground Truth Format: The ground truth data, which is essential for evaluating the accuracy of C-SLAM algorithms, is provided as TXT files. These files contain timestamped poses in UTM coordinates and orientation quaternions, formatted as follows: $\left[timestamp,t_{x},t_{y},t_{z},q_{x},q_{y},q_{z},q_{w}\right]$. The position $t_{a\in\{x,y,z\}}$ denotes the robot’s location in UTM coordinates, while $q_{a\in\{x,y,z,w\}}$ represents the orientation in quaternion form.
Sensor Data Topics: LABEL:Tab:message outlines the sensor topics included in the dataset, detailing the type of data each topic contains, its description, and the frequency of data capture.
IV Experiments
Table VI: ATE $[m]$ for Single SLAM and C-SLAM in the S3Ev1.0 outdoor environment without UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively. ✗ fails to initialize or track frames. If inter-loop closures detection fails, we mark it “Failed”.
Methods Square_1 Square_2 Library_1 Campus_Road_1 Playground_1 Playground_2 Dormitory_1
$\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$
ORB-SLAM3 1.16 15.5 - 2.81 11.4 55.5 0.87 3.29
VINS-Fusion 1.81 4.83 - 1.51 0.62 7.95 7.86 5.56 16.5 3.24 7.36 4.31 31.3 6.67 3.97 7.70
LIO-SAM 1.19 1.75 - 0.73 0.36 1.12 1.52 1.14 2.06 3.25 2.43 0.86 0.68 0.63 1.44 0.91
LVI-SAM 1.21 0.88 - 0.79 0.40 1.89 1.67 1.31 2.44 3.14 1.30 1.59 0.76 6.78 6.10 0.72 0.86 1.48 0.94
CoLRIO (front-end) 0.90 0.57 0.65 - 0.49 0.19 1.37 1.54 0.88 1.06 0.91 1.10 0.39 0.45 0.20 0.69 0.35 0.29 0.47 0.83 0.81
COVINS 1.63 0.83 1.75 Failed 18.3 9.61 54.7
DiSCo-SLAM Failed Failed 0.74 1.33 1.27 2.21 1.35 Failed 0.27 0.37 0.38 0.54 1.47 Failed
Swarm-SLAM 8.13 7.11 2.13 - 1.39 0.52 3.16 2.71 3.22 7.62 11.3 6.36 3.36 3.38 1.70 1.36 1.32 2.05 10.0 5.61 4.50
DCL-SLAM 1.23 0.86 0.66 - 0.50 0.26 0.58 1.26 1.17 1.51 1.48 1.87 0.33 0.40 0.32 0.52 1.37 0.51
CoLRIO 0.91 0.53 0.80 - 0.56 0.20 1.06 1.05 1.02 1.08 0.79 1.08 0.39 0.60 0.16 0.44 0.32 0.29 0.54 0.87 0.84
Table VII: ATE $[m]$ for Single SLAM and C-SLAM in the S3Ev1.0 indoor environment without UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively.
Methods Laboratory_1 Laboratory_2 Laboratory_3 Laboratory_4
$\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$
LIO-SAM 0.47 1.82 2.51 12.2 6.12 8.88 6.24 8.57 7.31 1.96 0.88 0.65
FAST-LIO2 7.27 11.2 5.13 10.2 2.14 15.0 12.8 6.80 7.33 0.79 0.88 0.37
CoLRIO (front-end) 0.40 0.51 0.44 0.83 0.79 0.96 1.79 0.91 2.43 0.67 0.65 0.87
DiSCo-SLAM 0.56 0.43 0.33 2.02 0.36 4.60 2.01 0.31 1.28 1.40 2.02 3.45
Swarm-SLAM 9.86 1.14 3.29 4.56 7.03 15.3 7.06 12.2 8.10 1.79 3.40 3.47
CoLRIO 0.43 0.50 0.44 0.40 0.07 0.08 0.49 0.91 0.13 0.95 0.65 0.18
Table VIII: ATE $[m]$ for C-SLAM in the S3Ev2.0 outdoor environment with UWB measurement. $\alpha$, $\beta$, and $\gamma$ stand for Alpha, Bob, and Carol respectively.
Methods Square_3 Library_2 Campus_Road_2 Campus_Road_3 Playground_3 Tunnel_1
$\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$ $\alpha$ $\beta$ $\gamma$
DiSCo-SLAM 1.15 1.26 12.11 1.07 1.02 0.98 15.41 14.60 3.37 1.27 1.32 39.76 9.01 0.72 2.32 5.06 11.90 4.40
Swarm-SLAM 8.52 5.64 2.80 3.15 0.91 0.94 48.96 64.23 22.37 8.79 9.60 4.90 1.11 3.24 5.47 10.49 11.06 10.06
CoLRIO (w/o UWB) 1.79 1.62 1.30 0.91 0.89 0.88 1.29 2.97 1.21 1.04 1.13 1.07 0.24 0.19 0.30 5.49 4.69 5.15
CoLRIO (w UWB) 1.71 1.52 1.54 0.89 0.87 0.88 1.37 1.59 1.34 0.77 0.61 0.90 0.23 0.20 0.30 4.01 3.21 3.60
Figure 7: Map in Laboratory_1 with CoRLIO.
(a) Tunnel_1 with DiSCo-SLAM.
(b) Tunnel_1 with Swarm-SLAM.
(c) Tunnel_1 with CoLRIO.
(d) Campus_Road_2 with DiSCo-SLAM.
(e) Campus_Road_2 with Swarm-SLAM.
(f) Campus_Road_2 with CoLRIO.
Figure 8: The qualitative results of outdoor environments.
IV-A Baselines
We have implemented four single-agent SLAM systems, namely ORB-SLAM3 [37], VINS-Fusion [38], LIO-SAM [39], and LVI-SAM [40]. Additionally, we have incorporated five collaborative SLAM (C-SLAM) systems into our study, which include COVINS [4], DiSCo-SLAM [5], Swarm-SLAM [29], DCL-SLAM [8], and CoLRIO [10]. These systems have been evaluated using the S3E dataset.
For most of the baselines, we only modify the intrinsic and extrinsic of the sensors and use the left camera for evaluation.
[…]
[autoplay,loop,width=]10videos/360/teaching_building/000096
(a) Teaching_building_1.
[…]
[autoplay,loop,width=]10videos/360/square_1/000197
(b) Square_1.
Figure 9: The animations of S3E dataset. For an optimal viewing experience, we recommend using Adobe Reader / KDE Okular , as these applications provide the necessary support for smooth playback and clear visualization of the animated sequences.
IV-B Results
Our experimental evaluation of the S3E dataset provides valuable insights into the performance of various state-of-the-art SLAM methodologies under diverse real-world conditions. The results, summarized in LABEL:Tab:baseline_1 and LABEL:Tab:baseline_2 , reveal the absolute trajectory error (ATE) for both single-agent and collaborative SLAM (C-SLAM) systems in outdoor and indoor environments without UWB measurement.
  • Key Findings:
    • Inter-Loop Closures Detection Impact: The results highlight the increased performance of C-SLAM systems when successful in detecting inter-loop closures, showcasing the benefits of collaborative information sharing among agents.
    • Front-end State Estimation Impact: The performance of C-SLAM is heavily influenced by the accuracy of its front-end state estimation. For example, CoLRIO’s success in certain scenarios is attributed to its effective front-end tracker.
  • Specific Observations:
    • In scenarios with large spatial overlap, C-SLAM systems leveraged inter-robot measurements to enhance state estimation accuracy. However, in areas with limited overlap, reducing drift remained a challenge.
    • The incorporation of UWB measurements in CoLRIO significantly improved localization robustness and accuracy, as demonstrated in LABEL:Tab:baseline_3, showcasing the benefits of additional ranging data for C-SLAM systems.
The results from our experiments on the S3E dataset underscore the importance of continuous innovation in C-SLAM algorithms. The varying performance across different scenarios and the impact of collaborative data sharing highlight the need for further research. The qualitative results presented in LABEL:Fig:qualitative_result_2 and LABEL:Fig:qualitative_results_3, along with the animations in LABEL:Fig:animations, visually complement the performance metrics, providing a deeper understanding of the C-SLAM systems’ behavior in diverse environments.
V Conclusion
In conclusion, the S3E dataset represents a significant advancement in the field of Collaborative Simultaneous Localization and Mapping (C-SLAM). Its expansive and multimodal nature, coupled with the introduction of four distinct trajectory paradigms, provides an unparalleled platform for evaluating C-SLAM systems across a multitude of real-world scenarios.Captured across various indoor and outdoor environments, the dataset includes meticulously synchronized streams from a range of sensors, offering an unprecedented level of detail and complexity. Our experiments utilizing this dataset have highlighted the improved robustness of C-SLAM systems, especially in their ability to handle inter-loop closures. The incorporation of UWB measurements has further enhanced the accuracy and reliability of localization within these systems. The S3E dataset serves not only as a benchmark for current C-SLAM methodologies but also as a catalyst for future innovations in multi-agent robotic navigation and mapping. We anticipate that this dataset will inspire further research and development, pushing the boundaries of what is possible in collaborative robotic systems.